# Deep Discriminative to Kernel Density Graph for In- and Out-of-distribution Calibrated Inference

### v7

@@ 5.2.2 Vision Benchmark Datasets Study
MCE $\downarrow$ $0.09\pm 0.00$ $\mathbf{0.04}\pm 0.00$ $\mathbf{0.03}\pm 0.01$ $0.10\pm 0.01$ $0.13\pm 0.00$ $0.09\pm 0.00$ $0.55\pm 0.00$
OOD CIFAR-100 AUROC $\uparrow$ $0.82\pm 0.00$ $\mathbf{0.83}\pm 0.00$ $\mathbf{0.83}\pm 0.00$ $0.80\pm 0.00$ $0.81\pm 0.00$ $0.81\pm 0.00$ $0.81\pm 0.00$
FPR@95 $\downarrow$ $\mathbf{0.50}\pm 0.33$ $\mathbf{0.50}\pm 0.26$ $\mathbf{0.54}\pm 0.26$ $\mathbf{0.51}\pm 0.32$ $\mathbf{0.56}\pm 0.27$ $\mathbf{0.49}\pm 0.33$ $\mathbf{0.45}\pm 0.35$
OCE $\downarrow$ $0.30\pm 0.00$ $0.20\pm 0.01$ $0.37\pm 0.01$ $0.29\pm 0.01$ $0.55\pm 0.00$ $0.31\pm 0.00$ $\mathbf{0.01}\pm 0.00$
SVHN AUROC $\uparrow$ $0.27\pm 0.00$ $\mathbf{1.00}\pm 0.00$ $0.22\pm 0.03$ $0.20\pm 0.07$ $0.36\pm 0.04$ $0.27\pm 0.01$ $0.70\pm 0.01$
FPR@95 $\downarrow$ $0.58\pm 0.42$ $\mathbf{0.14}\pm 0.11$ $0.63\pm 0.35$ $0.59\pm 0.41$ $0.64\pm 0.36$ $0.58\pm 0.43$ $0.54\pm 0.44$
OCE $\downarrow$ $0.87\pm 0.00$ $\textbf{0.01}\pm 0.00$ $0.85\pm 0.00$ $0.69\pm 0.01$ $0.90\pm 0.00$ $0.87\pm 0.00$ $0.04\pm 0.01$
Noise AUROC $\uparrow$ $0.20\pm 0.00$ $\mathbf{1.00}\pm 0.00$ $0.36\pm 0.07$ $0.24\pm 0.09$ $\mathbf{1.00}\pm 0.00$ $0.97\pm 0.00$ $0.97\pm 0.02$
FPR@95 $\downarrow$ $0.58\pm 0.42$ $\mathbf{0.12}\pm 0.12$ $0.64\pm 0.36$ $0.59\pm 0.41$ $0.38\pm 0.38$ $\mathbf{0.08}\pm 0.08$ $\mathbf{0.15}\pm 0.29$
OCE $\downarrow$ $0.90\pm 0.00$ $\mathbf{0.00}\pm 0.00$ $0.87\pm 0.00$ $0.71\pm 0.00$ $0.01\pm 0.01$ $0.06\pm 0.00$ $\mathbf{0.00}\pm 0.00$
For vision benchmark experiments, we use a Resnet-50 encoder pretrained using contrastive loss [27] as described in http://keras.io/examples/vision/supervised-contrastive-learning. The encoder projects the input images down to a $256$ dimensional latent space and we add two dense layers with $200$ and $10$ nodes on top of the encoder. We use the same pretrained encoder for all the baseline algorithms.
As shown in Table 1, KDN achieves good calibration for both ID and OOD datasets whereas the ID calibration approaches are poorly calibrated in the OOD regions and the OOD approaches have poor ID calibration.
6 Discussion
In this paper, we convert deep discriminative algorithms to kernel density graph by replacing the affine function over the polytopes in the discriminative networks with a Gaussian kernel. This replacement of affine function results in better in- and out-of-distribution calibration for our proposed approaches both on tabular and vision benchmark datasets. Our code, including the package and the experiments in this manuscript, will be made publicly available upon acceptance of the paper.
Acknowledgements
The authors thank the support of the NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning (NSF grant 2031985) and THEORINET. This work is graciously supported by the Defense Advanced Research Projects Agency (DARPA) Lifelong Learning Machines program through contracts FA8650-18-2-7834 and HR0011-18-2-0025. Research was partially supported by funding from Microsoft Research and the Kavli Neuroscience Discovery Institute.
Appendix A Proofs
A.1 Proof of Proposition 4.1
@@ Appendix A Proofs
For proving that $d$ is a valid distance metric for $\mathcal{P}_{n}$, we need to prove the following four statements:
@@
    Proof: Let $\mathcal{A}_{r}$ denote the set of activation modes in a ReLU-net and the set of leaf nodes in a random forest for a particular polytope $r$. $N$ is the total number of possible activation paths in a ReLU-net or total trees in a random forest. We can write:
    $N$ $\geq n((\mathcal{A}_{r}\cap\mathcal{A}_{s})\cup(\mathcal{A}_{s}\cap\mathcal{A}_{t}))$ (21)
    $=n(\mathcal{A}_{r}\cap\mathcal{A}_{s})+n(\mathcal{A}_{s}\cap\mathcal{A}_{t})-n(\mathcal{A}_{r}\cap\mathcal{A}_{s}\cap\mathcal{A}_{t})$
    $\geq n(\mathcal{A}_{r}\cap\mathcal{A}_{s})+n(\mathcal{A}_{s}\cap\mathcal{A}_{t})-n(\mathcal{A}_{r}\cap\mathcal{A}_{t})$
@@
    $N-n(\mathcal{A}_{r}\cap\mathcal{A}_{t})\leq N-n(\mathcal{A}_{r}\cap\mathcal{A}_{s})+N-n(\mathcal{A}_{s}\cap\mathcal{A}_{t})$
    $\implies 1-\frac{n(\mathcal{A}_{r}\cap\mathcal{A}_{t})}{N}\leq 1-\frac{n(\mathcal{A}_{r}\cap\mathcal{A}_{s})}{N}+1$
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\frac{n(\mathcal{A}_{s}\cap\mathcal{A}_{t})}{N}$
    $\implies d(r,t)\leq d(r,s)+d(s,t)$ (22)
A.2 Proof of Proposition 4.4
We first expand $\hat{g}_{y}(\mathbf{x})$:
@@
$\lim_{d_{\mathbf{x}}\to\infty}\tilde{f}_{y}(\mathbf{x})=0$
@@
$=\frac{(\lim_{d_{\mathbf{x}}\to\infty}\tilde{f}_{y}(\mathbf{x}))\hat{P}_{Y}(y)+\frac{b}{\log(n)}\hat{P}_{Y}(y)}{\sum_{k=1}^{K}(\lim_{d_{\mathbf{x}}\to\infty}\tilde{f}_{k}(\mathbf{x}))\hat{P}_{Y}(k)+\frac{b}{\log(n)}\hat{P}_{Y}(k))}$
@@
$=\hat{P}_{Y}(y)$
Appendix B Simulations
@@
Table 3: Hyperparameters for ReLU-net and KDN.
@@
Appendix C Pseudocodes
We provide the pseudocode for our porposed algorithms in Algorithm 1, 2 and 3.
Algorithm 1 Fit a KGX model.
1:
4:$\mathcal{G}$triangleright a KGX model
5:function KGX.fit($\theta,\mathbf{X},\mathbf{y}$)
7:         for $j=1,\ldots,n$ do
8:              $w_{ij}\leftarrow$ computeWeights($\mathbf{x}_{i},\mathbf{x}_{j},\theta$)
9:         end for
10:     end for
11:
12:
13:     $\{Q_{r}\}_{r=1}^{\tilde{p}}\leftarrow$ getPolytopes($\mathbf{w}$) $\triangleright$ Identify the polytopes by clustering the samples with similar weight
14:
15:     $\mathcal{G}.\{\tilde{w}_{k}\}_{k=1}^{K}\leftarrow 0$triangleright Initialize the counts for each class
16:     for $r=1,\ldots,\tilde{p}$ do $\triangleright$ Iterate over each polytope
17:         for $k=1,\ldots,K$ do
18:              $\mathcal{G}.\tilde{w}_{rk}\leftarrow$ countWeights($\{\mathbf{w}_{rs}\}_{s=1}^{\tilde{p}},k$) $\triangleright$ $w_{rk}$ is the number of weighted input samples in $Q_{r}$ with label $k$
19:              $\mathcal{G}.\tilde{w}_{k}\leftarrow\mathcal{G}.\tilde{w}_{k}+\mathcal{G}.\tilde{w}_{rk}$triangleright Update the total count for each class
20:         end for
21:         $\mathcal{G}.\hat{\mu}_{r},\mathcal{G}.\hat{\Sigma}_{r}\leftarrow$ estimateParameters($\mathbf{X},\{\mathbf{w}_{rs}\}_{s=1}^{\tilde{p}}$) $\triangleright$ Fit Gaussians using weighted MLE
22:     end for
23:     return $\mathcal{G}$
24:end function
Algorithm 2 Computing weights in KDF
1:
2:(1) $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbb{R}^{1\times d}$triangleright two input samples to be weighted
3:(2) $\theta$triangleright parent random forest with $T$ trees
4:$w_{ij}\in[0,1]$triangleright compute similarity between $i$ and $j$-th samples.
5:function computeWeights($\mathbf{x}_{i},\mathbf{x}_{j},\theta$)
6:     $\mathcal{I}_{i}\leftarrow$ pushDownTrees($\mathbf{x}_{i},\theta$) $\triangleright$ push $\mathbf{x}_{i}$ down $T$ trees and get the leaf numbers it end up in.
7:     $\mathcal{I}_{j}\leftarrow$ pushDownTrees($\mathbf{x}_{j},\theta$) $\triangleright$ push $\mathbf{x}_{j}$ down $T$ trees and get the leaf numbers it end up in.
8:     $l\leftarrow$ countMatches($\mathcal{I}_{i},\mathcal{I}_{j}$) $\triangleright$ count the number of times the samples end up in the same leaf
9:     $w_{ij}\leftarrow\frac{l}{T}$
10:     return $w_{ij}$
11:end function
Algorithm 3 Computing weights in KDN
1:
2:(1) $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbb{R}^{1\times d}$triangleright two input samples to be weighted
3:(2) $\theta$triangleright parent deep-net model
4:$w_{ij}\in[0,1]$triangleright compute similarity between $i$ and $j$-th samples.
5:function computeWeights($\mathbf{x}_{i},\mathbf{x}_{j},\theta$)
6:     $\mathcal{A}_{i}\leftarrow$ pushDownNetwork($\mathbf{x}_{i},\theta$) $\triangleright$ get activation modes $\mathcal{A}_{i}$
7:     $\mathcal{A}_{j}\leftarrow$ pushDownNetwork($\mathbf{x}_{j},\theta$) $\triangleright$ get activation modes $\mathcal{A}_{j}$
8:     $l\leftarrow$ countMatches($\mathcal{A}_{i}$, $\mathcal{A}_{j}$) $\triangleright$ count the number of times the two samples activate the activation paths in a similar way
9:     $w_{ij}\leftarrow\frac{l}{N}$triangleright $N$ is the total number of activation paths
10:     return $w_{ij}$
11:end function
Appendix D Time Complexity
[…]
Appendix E Hardware and Software Configurations
  • Operating System: Linux (ubuntu 20.04), macOS (Ventura 13.2.1)
  • VM Size: Azure Standard D96as v4 (96 vcpus, 384 GiB memory)
  • GPU: Apple M1 Max
Figure 5: Extended results on OpenML-CC18 datasets. Left: Performance (classification error, MCE and mean max confidence) of KDF on different Openml-CC18 datasets. Right: Performance (classification error, MCE and mean max confidence) of KDN on different Openml-CC18 datasets.
Figure 6: Extended results on OpenML-CC18 datasets (continued). Left: Performance (classification error, MCE and mean max confidence) of KDF on different Openml-CC18 datasets. Right: Performance (classification error, MCE and mean max confidence) of KDN on different Openml-CC18 datasets.
Figure 7: Extended results on OpenML-CC18 datasets (continued). Left: Performance (classification error, MCE and mean max confidence) of KDF on different Openml-CC18 datasets. Right: Performance (classification error, MCE and mean max confidence) of KDN on different Openml-CC18 datasets.
Figure 8: Extended results on OpenML-CC18 datasets (continued). Left: Performance (classification error, MCE and mean max confidence) of KDF on different Openml-CC18 datasets. Right: Performance (classification error, MCE and mean max confidence) of KDN on different Openml-CC18 datasets.

### v9

@@ Appendix A Frequently Asked Questions
OOD CIFAR-100 AUROC $\uparrow$ $0.88\pm 0.00$ $0.95\pm 0.01$ $\mathbf{0.96}\pm 0.00$ $0.92\pm 0.01$ $\mathbf{0.96}\pm 0.00$ $0.92\pm 0.01$ $0.92\pm 0.01$ $0.93\pm 0.00$ $0.93\pm 0.00$
FPR@95 $\downarrow$ $0.13\pm 0.01$ $0.12\pm 0.01$ $0.14\pm 0.01$ $0.14\pm 0.01$ $0.12\pm 0.00$ $0.34\pm 0.02$ $\mathbf{0.10}\pm 0.01$ $0.37\pm 0.01$ $0.39\pm 0.01$
SVHN AUROC $\uparrow$ $0.88\pm 0.01$ $0.97\pm 0.00$ $0.89\pm 0.00$ $0.90\pm 0.00$ $0.98\pm 0.00$ $0.94\pm 0.02$ $\mathbf{0.99}\pm 0.00$ $0.93\pm 0.02$ $0.94\pm 0.01$
FPR@95 $\downarrow$ $0.10\pm 0.02$ $0.08\pm 0.01$ $0.15\pm 0.00$ $0.12\pm 0.00$ $0.05\pm 0.01$ $0.17\pm 0.03$ $\mathbf{0.01}\pm 0.00$ $0.44\pm 0.09$ $0.45\pm 0.07$
Noise AUROC $\uparrow$ $0.85\pm 0.00$ $\mathbf{0.98}\pm 0.01$ $0.82\pm 0.00$ $0.89\pm 0.00$ $\mathbf{0.99}\pm 0.00$ $\mathbf{0.99}\pm 0.00$ $\mathbf{0.99}\pm 0.01$ $0.89\pm 0.06$ $0.86\pm 0.03$
FPR@95 $\downarrow$ $0.17\pm 0.00$ $0.03\pm 0.03$ $0.12\pm 0.00$ $0.11\pm 0.01$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.00}\pm 0.01$ $0.78\pm 0.31$ $0.99\pm 0.00$
[…]
OOD CIFAR-10 AUROC $\uparrow$ $0.89\pm 0.01$ $\mathbf{0.97}\pm 0.00$ $0.83\pm 0.02$ $0.90\pm 0.01$ $0.88\pm 0.02$ $0.90\pm 0.01$ $0.90\pm 0.01$ $0.81\pm 0.00$ $0.81\pm 0.01$
FPR@95 $\downarrow$ $0.41\pm 0.02$ $\mathbf{0.05}\pm 0.01$ $0.61\pm 0.04$ $0.38\pm 0.03$ $0.57\pm 0.04$ $0.58\pm 0.02$ $0.50\pm 0.03$ $0.66\pm 0.01$ $0.68\pm 0.01$
SVHN AUROC $\uparrow$ $0.91\pm 0.01$ $\mathbf{0.98}\pm 0.00$ $0.87\pm 0.01$ $0.91\pm 0.01$ $0.88\pm 0.02$ $0.81\pm 0.02$ $0.92\pm 0.01$ $0.0.89\pm 0.02$ $0.89\pm 0.02$
FPR@95 $\downarrow$ $0.39\pm 0.03$ $\mathbf{0.04}\pm 0.01$ $0.59\pm 0.00$ $0.38\pm 0.03$ $0.51\pm 0.06$ $0.62\pm 0.05$ $0.34\pm 0.03$ $0.89\pm 0.02$ $0.51\pm 0.06$
Noise AUROC $\uparrow$ $0.98\pm 0.01$ $\mathbf{1.00}\pm 0.00$ $0.93\pm 0.01$ $0.98\pm 0.01$ $\mathbf{1.00}\pm 0.00$ $0.98\pm 0.01$ $0.98\pm 0.02$ $0.99\pm 0.01$ $\mathbf{1.00}\pm 0.00$
FPR@95 $\downarrow$ $0.06\pm 0.05$ $\mathbf{0.00}\pm 0.00$ $0.42\pm 0.03$ $0.07\pm 0.06$ $0.01\pm 0.00$ $0.08\pm 0.04$ $0.15\pm 0.17$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.00}\pm 0.00$
[…]
OOD CIFAR-10 AUROC $\uparrow$ $0.95\pm 0.01$ $0.97\pm 0.01$ $0.96\pm 0.01$ $0.94\pm 0.01$ $0.95\pm 0.01$ $0.96\pm 0.01$ $\mathbf{0.99}\pm 0.00$ $0.52\pm 0.01$ $0.54\pm 0.01$
FPR@95 $\downarrow$ $0.30\pm 0.07$ $0.06\pm 0.02$ $0.25\pm 0.08$ $0.34\pm 0.09$ $0.38\pm 0.06$ $0.18\pm 0.07$ $\mathbf{0.02}\pm 0.01$ $0.97\pm 0.00$ $0.97\pm 0.00$
CIFAR-100 AUROC $\uparrow$ $0.95\pm 0.01$ $\mathbf{0.97}\pm 0.01$ $0.96\pm 0.01$ $0.95\pm 0.01$ $0.95\pm 0.01$ $0.95\pm 0.01$ $0.95\pm 0.01$ $0.52\pm 0.01$ $0.55\pm 0.00$
FPR@95 $\downarrow$ $0.30\pm 0.07$ $\mathbf{0.06}\pm 0.02$ $0.25\pm 0.07$ $0.35\pm 0.08$ $0.37\pm 0.05$ $0.19\pm 0.06$ $\mathbf{0.04}\pm 0.02$ $0.97\pm 0.00$ $0.97\pm 0.00$
Noise AUROC $\uparrow$ $0.96\pm 0.02$ $0.98\pm 0.03$ $0.98\pm 0.01$ $0.96\pm 0.01$ $\mathbf{1.00}\pm 0.00$ $0.99\pm 0.00$ $\mathbf{1.00}\pm 0.00$ $0.63\pm 0.06$ $0.64\pm 0.06$
FPR@95 $\downarrow$ $0.24\pm 0.16$ $0.04\pm 0.05$ $0.16\pm 0.11$ $0.26\pm 0.13$ $\mathbf{0.00}\pm 0.00$ $0.01\pm 0.01$ $0.01\pm 0.01$ $1.0\pm 0.00$ $1.0\pm 0.00$
Why did you choose CIFAR-10, CIFAR-100 and SVHN for your vision experiments?
We have done vision experiments using CIFAR-10 (10 classes), CIFAR-100 (100 classes) and SVHN (10 classes and bigger training size) as ID datasets. We emphasize that CIFAR-10, CIFAR-100 and SVHN are some of the hardest ID and OOD pairs according to various papers (Nalisnick et al., 2018; Fort et al., 2021), and hence they are adopted as the benchmarking datasets by many of the papers in the literature. Doing experiments with extremely large datasets like imagenet (14 million images) is computationally and storage-wise expensive using our current implementation. We will pursue extremely large datasets in future. Many relevant papers on OOD calibration use only small- and mid-sized datasets (Gardner et al., 2024; Borisov et al., 2022; Ulmer et al., 2020). We acknowledge this limitation in the paper.
As shown in Algorithm 2, in each iteration, Geodesic kernel values between the current training input and the entire dataset need to be computed, each of which involves comparing all possible activation paths through the neural network (Equation 12 and Algorithm 1). And Algorithm 2 iterates over every single training instance. This will become extremely challenging on modern neural networks with miliions, or even billions, of parameters (nodes) trained on large-scale datasets consisting of millions of training samples.
We address this scalability issue with two solutions:
  1. 1.
    Parallelization for Efficient Kernel Computation: Our kernel is computed as the product of Hamming distances between activation patterns (binary strings of ‘1’ and ‘0’) at each layer for two samples. Since these calculations are independent across layers and sample pairs, they can be efficiently parallelized. We leverage the cdist function from scipy.spatial.distance, which enables fast pairwise distance computations. This parallel implementation reduced runtime from days to just minutes.
  2. 2.
    Efficient Computation for Large Models: For extremely large models, such as vision transformers (ViTs) in Section 3.2.2, we employ a front-end encoder to extract local image features (as described in Section 3.2, first paragraph). Instead of applying KDN to the entire ViT, we fit it only to the final fully connected layers. This reduces the size of the model by about $90\%$ while maintaining performance, allowing more efficient KDN computation.
As dimension continues to increase, densities of a Gaussian kernel tend to concentrate on a thin shell, where densities decay exponentially both going outward (away from the centre) and inward (towards the centre). Thus, a low class conditional density estimated by the Gaussian kernel does not necessarily imply the inference sample is OOD. Such high-dimensional feature space (large ) is commonly encountered in practical tasks.
In such high dimensional spaces, most of the probability mass concentrates on a thin shell at a certain radius from the mean due to the rapid growth of volume on the surface of the shell. While the density of a Gaussian distribution is highest at the mean and decays exponentially with Euclidean distance from the mean, the total probability mass at a given radius is given by the integral of the probability density over that region. The volume of the surface at that radius increases rapidly with higher dimensions (Fraser, 1984). In high dimensions, this results in the counterintuitive effect where the peak probability mass is not at the mean but at some distance from it. However, this phenomenon does not affect our approach. As shown in Equation 5, we evaluate the Gaussian kernel at a single point x, not over an entire surface or volume. Since the kernel density is maximized at the center and decays exponentially outward, the volume concentration effect does not alter the inference process in our method.
Appendix B Proofs
B.1 Proof of Proposition 1
@@
To prove that $d$ is a valid distance metric for $\mathcal{P}_{n}$, we need to prove the following four statements. Here we use the notation $d(r,s)$ to denote the distance between two samples $\boldsymbol{x}_{i}\in Q_{r}$ and $\boldsymbol{x}_{s}\in Q_{s}$.
@@
    Proof: Let $\mathcal{A}_{r}$ denote the sequence of elements in the activation mode for a particular polytope $r$. $B$ is the cardinality of $\mathcal{A}_{r}$. Below $c(\cdot)$ denotes the cardinality of the sequence. We can write:
    $B$ $\geq c((\mathcal{A}_{r}\cap\mathcal{A}_{s})\cup(\mathcal{A}_{s}\cap\mathcal{A}_{t}))$ (26)
    $=c(\mathcal{A}_{r}\cap\mathcal{A}_{s})+c(\mathcal{A}_{s}\cap\mathcal{A}_{t})-c(\mathcal{A}_{r}\cap\mathcal{A}_{s}\cap\mathcal{A}_{t})$
    $\geq c(\mathcal{A}_{r}\cap\mathcal{A}_{s})+c(\mathcal{A}_{s}\cap\mathcal{A}_{t})-c(\mathcal{A}_{r}\cap\mathcal{A}_{t}).$
@@
    $B-c(\mathcal{A}_{r}\cap\mathcal{A}_{t})\leq B-c(\mathcal{A}_{r}\cap\mathcal{A}_{s})+B-c(\mathcal{A}_{s}\cap\mathcal{A}_{t})$
    $\implies 1-\frac{c(\mathcal{A}_{r}\cap\mathcal{A}_{t})}{B}\leq 1-\frac{c(\mathcal{A}_{r}\cap\mathcal{A}_{s})}{B}+1$
    $~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\frac{c(\mathcal{A}_{s}\cap\mathcal{A}_{t})}{B}$
    $\implies d(r,t)\leq d(r,s)+d(s,t).$ (27)
B.2 Proof of Proposition 2
Note that first we find the nearest polytope to the inference point $x$ using Geodesic distance and use Gaussian kernel locally for $x$ within that polytope. Here the Gaussian kernel uses Euclidean distance from the kernel center to $x$ (within the numerator of the exponent). The value out of the Gaussian kernel decays exponentially with the increasing distance of the inference point from the kernel center. We first expand $\hat{g}_{y}(\mathbf{x})$:
@@ B.2 Proof of Proposition 2
$\lim_{d_{\mathbf{x}}\to\infty}\tilde{f}_{y}(\mathbf{x})=0.$
@@
$=\hat{P}_{Y}(y).$
Note that without the constant term $\frac{b}{\log(n)}$, the class priors in the numerator will be 00 and hence we will not have Proposition 2.
Appendix C Hardware and Software Configurations
  • Operating System: Linux (ubuntu 20.04), macOS (Ventura 13.2.1)
  • VM Size: Azure Standard D96as v4 (96 vcpus, 384 GiB memory)
  • GPU: Apple M1 Max
Appendix D Simulations
Figure 7: Additional simulation result. KDF and KDN improve both in- and out-of-distribution calibration of their respective parent algorithms while maintaining nearly similar classification accuracy on the simulation datasets.
@@ Appendix D Simulations
$b$ $\exp{(-10^{-7})}$
Table 4: Hyperparameters for ReLU-net and KDN on Tabular data.
@@
$b$ $\exp{(-10^{-7})}$
Appendix E Extended Results on OpenML-CC18 data suite
See Figure 9, 10, 11 and 12 for extended results on OpenML-CC18 data suite.
Figure 9: Extended results on OpenML-CC18 datasets. Left: Performance (classification error, ECE and mean max confidence) of KDF on different Openml-CC18 datasets. Right: Performance (classification error, ECE and mean max confidence) of KDN on different Openml-CC18 datasets.
Figure 10: Extended results on OpenML-CC18 datasets (continued). Left: Performance (classification error, ECE and mean max confidence) of KDF on different Openml-CC18 datasets. Right: Performance (classification error, ECE and mean max confidence) of KDN on different Openml-CC18 datasets.
Figure 11: Extended results on OpenML-CC18 datasets (continued). Left: Performance (classification error, ECE and mean max confidence) of KDF on different Openml-CC18 datasets. Right: Performance (classification error, ECE and mean max confidence) of KDN on different Openml-CC18 datasets.
Figure 12: Extended results on OpenML-CC18 datasets (continued). Left: Performance (classification error, ECE and mean max confidence) of KDF on different Openml-CC18 datasets. Right: Performance (classification error, ECE and mean max confidence) of KDN on different Openml-CC18 datasets.